# FutureSightDrive: Thinking Visually with Spatio-Temporal CoT for Autonomous Driving

### v2

@@ Abstract
Vision–Language–Action (VLA) models are increasingly used for end-to-end driving due to their world knowledge and reasoning ability. Most prior work, however, inserts textual chains-of-thought (CoT) as intermediate steps tailored to the current scene. Such symbolic compressions can blur spatio-temporal relations and discard fine visual cues, creating a cross-modal gap between perception and planning. We propose FSDrive, a visual spatio-temporal CoT framework that enables VLAs to think in images. The model first acts as a world model to generate a unified future frame that overlays coarse but physically-plausible priors—future lane dividers and 3D boxes—on the predicted future image. This unified frame serves as the visual CoT, capturing both spatial structure and temporal evolution. The same VLA then functions as an inverse-dynamics model, planning trajectories from current observations and the visual CoT. To equip VLAs with image generation while preserving understanding, we introduce a unified pre-training paradigm that expands the vocabulary to include visual tokens and jointly optimizes VQA (for semantics) and future-frame prediction (for dynamics). A progressive easy-to-hard scheme first predicts lane/box priors to enforce physical constraints, then completes full future frames for fine details. On nuScenes and NAVSIM, FSDrive improves trajectory accuracy and reduces collisions under both ST-P3 and UniAD metrics, and attains competitive FID for future-frame generation despite using lightweight autoregression. It also advances scene understanding on DriveLM. Together, these results indicate that visual CoT narrows the cross-modal gap and yields safer, more anticipatory planning. Code is available at https://github.com/MIV-XJTU/FSDrive.
@@ 1 Introduction
Recently, given the superior capabilities of multimodal large language models (MLLMs) in world knowledge, reasoning ability, and interpretability, they have been widely applied in autonomous driving [20, 43, 87, 31]. One promising direction is the end-to-end vision-language-action (VLA) model, which leverages pre-trained vision-language model (VLM) to directly extract scene features from visual observations and language instructions, subsequently generating vehicle control commands (e.g., speed and trajectory). This paradigm not only simplifies system architecture and minimizes information loss, but also enables the utilization of the model’s world knowledge to analyze driving environments and reason about safe decisions in complex scenarios.
In the field of language, Chain-of-Thought (CoT) [67, 50, 15, 52] improves reasoning capabilities and interpretability by encouraging step-by-step thinking. However, existing autonomous driving [27, 44, 14] typically incorporate discrete text CoT (e.g., language descriptions targeting current scenarios and bounding box coordinates) as intermediate reasoning steps. This method is essentially highly abstract and symbolized compression of visual information, which may lead to ambiguous temporal and spatial relationships, loss of fine-grained information, and modality conversion gaps [46, 55, 72], as shown in the top of Figure 1. For autonomous vehicles requiring deep physical-world interaction, should their thinking process more closely resemble simulation and imagination of world, rather than merely relying on logical deduction of language?
@@
Inspired by the human driver’s cognitive mechanism of directly constructing visual representations of future scenarios in the mind, rather than converting them into language descriptions for reasoning, we propose a more intuitive spatio-temporal CoT method as shown in the bottom part of Figure 1. This method avoids information loss during text abstraction and enables the model to think visually about trajectory planning. First, the VLA serves as a world model to generate unified image frame for predicting future world states: Inspired by visual prompting engineering [53, 81] that draws red circles on images to guide model attention and by VLIPP [78] first predicts future bounding boxes to introduce physical priors when generating future frames, we represent future world spatial relationships through future red lane dividers and 3D detection boxes on the predicted unified frames [80]. These coarse-grained visual cues direct the model’s attention toward drivable areas and critical objects in future scenes while enforcing physically plausible constraints. Meanwhile, the temporal relationships are represented by the ordinary future frame, where the dynamic evolution of visual content intuitively characterizes temporal progression and the inherent laws of scene development. Subsequently, the spatio-temporal CoT acts as an intermediate reasoning step, enabling the VLA to function as an inverse dynamics model for trajectory planning based on current observations and future predictions. Compared to traditional discrete text CoT, and even image-text CoT methods [27, 91, 41] as shown in the middle of the Figure 1, our method unifies both future scene representations and perception outputs in image format, which more effectively conveys the temporal and spatial relationships. This eliminates semantic gaps caused by cross-modal conversions (e.g., converting visual perceptions into textual descriptions for reasoning), establishing an end-to-end visual reasoning pipeline that enables direct visual causal inference by the model.
@@
To endow VLAs with image generation capabilities, we propose a pre-training paradigm that simultaneously preserves the semantic understanding of existing MLLM and activates their visual generation capacity. Specifically, for the semantic understanding preservation part, we follow previous approaches [64, 27, 25] by incorporating visual question answering (VQA) tasks for current scene comprehension. For the activation of visual generation capabilities, we investigate the shared vocabulary space between image and text, directly unleashing the visual generation potential of existing MLLMs in the field of autonomous driving through minimal data (approximately 0.3% of previous methods [70, 73, 24, 35]) without requiring complex model architecture modifications or redesigns. However, directly generating complete detailed future scenes may fail to adhere to physical laws [78, 88]. Thus, we propose a progressive, easy-to-hard generation method. We leverage the world knowledge of VLAs to first infer drivable regions and key object positions in future scenarios, generating coarse-grained future perception images (e.g., lane dividers and 3D detection) to constrain physical laws. Subsequently, full future frames are generated under this constraint to supplement fine-grained details, enabling the model to think visually about accurate future prediction.
@@ 2.1 Unified multimodal understanding and generation
Recent research efforts [38, 70, 49, 68] have increasingly focused on unifying multimodal understanding and visual generation within a single LLM. On one front, methods like Show-o [74], and VILA-U [73] employ VQ-VAE [61] to transform images into discrete tokens while training LLMs to predict them. However, these methods suffer from insufficient semantic information preservation, often leading to performance degradation in downstream understanding tasks. Alternative methods [57, 11, 48, 9, 82] utilize ViT [12]-based vision encoders (e.g., CLIP [51]) to encode images into continuous feature maps. Nevertheless, such methods typically depend on external diffusion models for image generation or use different training objectives (i.e. diffusion and autoregression) for the two tasks, further complicates the infrastructure design with overall lower efficiency. Moreover, the aforementioned methods usually require massive billion-scale datasets for extensive training from scratch, which results in prohibitively high computational costs when disseminating explorations in this form. In this work, we demonstrate that the visual generative capabilities of existing MLLMs can be directly activated through minimal training costs (approximately 0.3% of previous methods [70, 58, 42, 8]) without requiring sophisticated architectural designs.
@@ 2.2 Vision-language models for autonomous driving
Given the superior capabilities of large language models (LLMs) in world knowledge, reasoning, and interpretability, recent researches [2, 83, 39, 85] increasingly integrate Vision-Language Models (VLMs)/LLMs with autonomous driving systems to address limitations in end-to-end approaches. DriveGPT4 [76] employs LLMs through iterative question-answering interactions to explain vehicle behaviors and predict control signals. DriveVLM [60] synergizes LLMs with end-to-end architectures, where LLMs predict low-frequency trajectories that are subsequently refined by the end-to-end model for final planning. Doe-1 [95] reformulates autonomous driving as a next-token prediction task using Lumina-mGPT’s [37] multimodal generation capabilities, executing diverse tasks through multimodal token processing. EMMA [27] leverages Gemini’s multimodal foundation by encoding all non-sensor inputs (navigation instructions, vehicle status) and outputs (trajectories, 3D positions) as natural language text, fully exploiting pre-trained LLMs’ world knowledge. In this work, we propose a spatio-temporal chain of thought (CoT) reasoning method that unifies the form of images, allowing the model to think visually about trajectory planning.
@@ 2.3 World models for autonomous driving
World models [66, 45, 90, 89] aim to infer ego status and dynamic environments from past observations to enable accurate future prediction and planning. Current applications of world models in autonomous driving primarily focus on driving scenario generation [47, 16, 32], planning [66, 41], and representation learning [45, 79, 84]. For driving scenario generation, most prior works are built upon diffusion models, with the exception of GAIA-1 [18] which incorporates a progressive next-token predictor and an additional diffusion image decoder. Recent DrivingGPT [5] leverages existing vision generation LLM LlamaGen [56] while simultaneously outputting predictions for future states and actions. However, such VQ-VAE based visual tokens lack semantic information, often leading to performance degradation in downstream visual understanding tasks [74, 40, 59]. In this work, we propose to directly activate the visual generation capabilities of existing multimodal large language models, enabling VLMs to act as world models and predict future frames.
@@ 3.1 Preliminary
End-to-end autonomous driving directly generates future trajectory from sensor data, convertible to vehicle control actions like acceleration and steering [27]. given $N$ surround-view images $I_{t}=\{I_{t}^{1},I_{t}^{2},\ldots,I_{t}^{N}\}$ at timestep $t$, model $\mathcal{M}$ outputs a BEV trajectory $W_{t}=\{w_{t}^{1},w_{t}^{2},\ldots,w_{t}^{n}\}$, where each waypoint $w_{t}^{i}=(x_{t}^{i},y_{t}^{i})$. The process is formulated as:
@@
Recent works [70, 22] unify multimodal understanding and vision generation in single LLM. While understanding aligns with standard LLMs, generation methods [38, 23] typically use VQ-VAE [61] to encode images into discrete tokens. First, the image tokenizer quantizes image pixels $x\in\mathbb{R}^{H\times W\times 3}$ into discrete tokens $q\in\mathcal{Q}^{h\times w}$, where $h=H/p$, $w=W/p$, $p$ is the downsampling factor, and $q(i,j)$ represents the index of the image codebook. These $h\cdot w$ tokens are arranged in raster order to train a Transformer [62]-based autoregressive model. During image generation, a general language modeling (LM) objective is adopted to autoregressively predict the next token, maximizing the likelihood of each image token:
@@ 3.2 Unified pre-training paradigm for visual generation and understanding
To enable unified pre-training, MLLMs require visual generation capabilities. As described in Section 3.1, existing methods (e.g. Lumina-mGPT [37], the visual generation LLM used by Doe-1 [95]) typically employ VQ-VAE to encode images into discrete tokens when extracting visual information. However, these tokens lack semantic information, which hurts downstream understanding performance [74, 97]. Moreover, current methods [70, 96] demand expensive training from scratch on massive billion-scale datasets without leveraging existing LLM knowledge.
@@
To effectively preserve the semantic understanding capabilities of the native MLLM during the pre-training stage, as shown in the left part of Figure 2, we follow previous methods [64, 27] by using a VQA task, which is crucial for autonomous vehicles to analyze complex driving scenarios. Given an image-text question-answer pair $(I,L)$, where $I$ represents the surround-view images of the current scene and $L$ denotes the instructional question, model $\mathcal{M}$ generates a corresponding answer $A$:
@@
Inspired by the world models in autonomous driving [30, 77] that generate future frames to learn physical laws, after activating the visual generation capability, we also enable the VLA to predict future frames, thereby capturing the dynamic evolution of the world. Specifically, given an image-instruction pair $(I,L)$, the model predicts the next visual token of the future front-view frame through autoregressive generation:
@@
However, directly generating complete detailed future scenes may fail to adhere to physical laws [78]. Therefore, during pre-training stage, we propose a progressive, easy-to-hard generation method, incorporating annotated data containing lane divider and 3D detection. Before generating visual tokens of future frames $Q_{f}$, we leverage the world knowledge of VLA to first reason about visual tokens of lane dividers $Q_{l}$, which serve as the skeleton of the road scene and define drivable areas to enforce static physical constraints. Subsequently, we reason about visual tokens of 3D bounding boxes $Q_{d}$, representing motion patterns of key objects to impose dynamic physical constraints. This progressive method sequence explicitly guides the model to infer structural layouts and geometric details of future scenes while enforcing physical laws. By leveraging these intermediate visual reasoning steps as context, the model learns to think visually about the dynamic evolution of scenes, ultimately enabling accurate future prediction:
@@ 3.3 Think visually with spatio-temporal CoT
Autonomous driving planning requires not only understanding the current scene but also envisioning potential future developments to achieve forward-looking comprehension. This thinking process should resemble physical world simulation and imagination rather than purely text symbolic logical deduction. Since our model has already learned physical constraints through the progressive generation during pre-training, and considering efficiency, we no longer separately generate lane dividers, 3D detection, and future frames, but instead integrate all these results into a single unified frame. As shown in the right part of Figure 2, here, VLA serves as a world model to generate a unified image frame predicting the future world state: Inspired by visual prompting engineering [53] that draws red circles on images to guide model attention and by VLIPP [78] first predicts future bounding boxes to introduce physical priors when generating future frames, we represent future world spatial relationships through future red lane dividers and 3D detection boxes on the predicted unified frames. These coarse-grained visual cues direct the model’s attention toward drivable areas and critical objects in future scenes while enforcing physically plausible constraints. Meanwhile, the temporal relationships are represented by the ordinary future frame, where the dynamic evolution of visual content intuitively characterizes temporal progression and the inherent laws of scene development. Subsequently, spatio-temporal CoT $Q_{CoT}$ serves as an intermediate reasoning step, allowing the VLA to function as an inverse dynamics model that plans trajectory based on current observations and future predictions:

### v3

@@ Abstract
Vision-Language-Action (VLA) models offer significant potential for end-to-end driving, yet their reasoning is often constrained by textual Chains-of-Thought (CoT). This symbolic compression of visual information creates a modality gap between perception and planning by blurring spatio-temporal relations and discarding fine-grained cues. We introduce FSDrive, a framework that empowers VLAs to "think visually" using a novel visual spatio-temporal CoT. FSDrive first operates as a world model, generating a unified future frame that combines a predicted background with explicit, physically-plausible priors like future lane dividers and 3D object boxes. This imagined scene serves as the visual spatio-temporal CoT, capturing both spatial structure and temporal evolution in a single representation. The same VLA then functions as an inverse-dynamics model to plan trajectories conditioned on current observations and this visual CoT. We enable this with a unified pre-training paradigm that expands the model’s vocabulary with visual tokens and jointly optimizes for semantic understanding (VQA) and future-frame prediction. A progressive curriculum first generates structural priors to enforce physical laws before rendering the full scene. Evaluations on nuScenes and NAVSIM show FSDrive improves trajectory accuracy and reduces collisions, while also achieving competitive FID for video generation with a lightweight autoregressive model and advancing scene understanding on DriveLM. These results confirm that our visual spatio-temporal CoT bridges the perception-planning gap, enabling safer, more anticipatory autonomous driving. Code is available at https://github.com/MIV-XJTU/FSDrive.
@@ 1 Introduction
The advent of Multimodal Large Language Models (MLLMs) is reshaping autonomous driving, with Vision-Language-Action (VLA) models emerging as a promising end-to-end paradigm [hu2025contextalignment, ma2023dolphins, zhang2024chatscene, li2024llada]. Harnessing the superior capabilities of MLLMs in world knowledge, reasoning, and interpretability, these models directly map visual observations and language instructions to vehicle control commands (e.g., speed and trajectory). This approach not only simplifies the conventional modular architecture, thereby minimizing potential information loss across components, but also enables the system to leverage vast pre-trained knowledge for analyzing complex driving environments and reasoning about safe decisions.
To enhance their reasoning abilities, many such models have incorporated the Chain-of-Thought (CoT) strategy, which encourages step-by-step thinking [Wei2022ChainOT, qu2025spatialvla, guo2025tom, sarkar2025reasoning]. However, in existing autonomous driving applications [Hwang2024EMMAEM, mao2023agentdriver, gao2025openfly], this often involves generating discrete textual CoTs (e.g., language descriptions of the current scene or bounding box coordinates) as intermediate steps. This process forces a conversion of rich, continuous visual data into abstract, symbolic representations — a form of lossy compression that can obscure critical spatio-temporal relationships, discard fine-grained visual details, and introduce a "modality gap" between perception and planning [motamed2025physics, song2025hume, wu2025teaching], as illustrated in Figure 1. For autonomous vehicles requiring deep physical-world interaction, should their thinking process more closely resemble simulation and imagination of world, rather than merely relying on logical deduction of language?
@@
Inspired by the human driver’s cognitive mechanism of directly constructing visual representations of future scenarios in the mind, rather than converting them into language descriptions for reasoning, we propose a more intuitive spatio-temporal CoT method as shown in the bottom part of Figure 1. This method avoids information loss during text abstraction and enables the model to think visually about trajectory planning. First, the VLA serves as a world model to generate unified image frame for predicting future world states: Inspired by visual prompting engineering [redcircle, yu2025forgetme] that draws red circles on images to guide model attention and by VLIPP [Yang2025VLIPPTP] first predicts future bounding boxes to introduce physical priors when generating future frames, we represent future world spatial relationships through future red lane dividers and 3D detection boxes on the predicted unified frames [yu2025physics]. These coarse-grained visual cues direct the model’s attention toward drivable areas and critical objects in future scenes while enforcing physically plausible constraints. Meanwhile, the temporal relationships are represented by the ordinary future frame, where the dynamic evolution of visual content intuitively characterizes temporal progression and the inherent laws of scene development. Subsequently, the spatio-temporal CoT acts as an intermediate reasoning step, enabling the VLA to function as an inverse dynamics model for trajectory planning based on current observations and future predictions. Compared to traditional discrete text CoT, and even image-text CoT methods [Hwang2024EMMAEM, Zhao2025CoTVLAVC, liu2025qfft] as shown in the middle of the Figure 1, our method unifies both future scene representations and perception outputs in image format, which more effectively conveys the temporal and spatial relationships. This eliminates semantic gaps caused by cross-modal conversions (e.g., converting visual perceptions into textual descriptions for reasoning), establishing an end-to-end visual reasoning pipeline that enables direct visual causal inference by the model.
@@
To endow VLAs with image generation capabilities, we propose a pre-training paradigm that simultaneously preserves the semantic understanding of existing MLLM and activates their visual generation capacity. Specifically, for the semantic understanding preservation part, we follow previous approaches [Wang2024OmniDriveAH, Hwang2024EMMAEM, huang2025intelligent] by incorporating visual question answering (VQA) tasks for current scene comprehension. For the activation of visual generation capabilities, we investigate the shared vocabulary space between image and text, directly unleashing the visual generation potential of existing MLLMs in the field of autonomous driving through minimal data (approximately 0.3% of previous methods [wu2024janus, wu2024vila-u, huang2024intelligent, liang2025persistent]) without requiring complex model architecture modifications or redesigns. However, directly generating complete detailed future scenes may fail to adhere to physical laws [Yang2025VLIPPTP, zhang2025rearank]. Thus, we propose a progressive, easy-to-hard generation method. We leverage the world knowledge of VLAs to first infer drivable regions and key object positions in future scenarios, generating coarse-grained future perception images (e.g., lane dividers and 3D detection) to constrain physical laws. Subsequently, full future frames are generated under this constraint to supplement fine-grained details, enabling the model to think visually about accurate future prediction.
@@ 2.1 Unified multimodal understanding and generation
Recent research efforts [lwm, wu2024janus, Siamese-Diffusion, wei2025copeft] have increasingly focused on unifying multimodal understanding and visual generation within a single LLM. On one front, methods like Show-o [xie2024showo], and VILA-U [wu2024vila-u] employ VQ-VAE [vqvae] to transform images into discrete tokens while training LLMs to predict them. However, these methods suffer from insufficient semantic information preservation, often leading to performance degradation in downstream understanding tasks. Alternative methods [emu, dong2024dreamllm, qian2025diffusion, dai2025caption, yuan2025unimapgen] utilize ViT [vit]-based vision encoders (e.g., CLIP [clip]) to encode images into continuous feature maps. Nevertheless, such methods typically depend on external diffusion models for image generation or use different training objectives (i.e. diffusion and autoregression) for the two tasks, further complicates the infrastructure design with overall lower efficiency. Moreover, the aforementioned methods usually require massive billion-scale datasets for extensive training from scratch, which results in prohibitively high computational costs when disseminating explorations in this form. In this work, we demonstrate that the visual generative capabilities of existing MLLMs can be directly activated through minimal training costs (approximately 0.3% of previous methods [wu2024janus, sun2025gdiffretro, lu2025dammfnd, dai2025securetugo]) without requiring sophisticated architectural designs.
@@ 2.2 Vision-language models for autonomous driving
Given the superior capabilities of large language models (LLMs) in world knowledge, reasoning, and interpretability, recent researches [Chang2024MapDR, yue2025real, liu2024fedbcgd, zeng2025janusvln] increasingly integrate Vision-Language Models (VLMs)/LLMs with autonomous driving systems to address limitations in end-to-end approaches. DriveGPT4 [xu2024drivegpt4] employs LLMs through iterative question-answering interactions to explain vehicle behaviors and predict control signals. DriveVLM [DriveVLM] synergizes LLMs with end-to-end architectures, where LLMs predict low-frequency trajectories that are subsequently refined by the end-to-end model for final planning. Doe-1 [doe] reformulates autonomous driving as a next-token prediction task using Lumina-mGPT’s [liu2024lumina-mgpt] multimodal generation capabilities, executing diverse tasks through multimodal token processing. EMMA [Hwang2024EMMAEM] leverages Gemini’s multimodal foundation by encoding all non-sensor inputs (navigation instructions, vehicle status) and outputs (trajectories, 3D positions) as natural language text, fully exploiting pre-trained LLMs’ world knowledge. In this work, we propose a spatio-temporal chain of thought (CoT) reasoning method that unifies the form of images, allowing the model to think visually about trajectory planning.
@@ 2.3 World models for autonomous driving
World models [wang2023driving, Min2024DriveWorld4P, zhang2025cchall, zhang2025vitcot] aim to infer ego status and dynamic environments from past observations to enable accurate future prediction and planning. Current applications of world models in autonomous driving primarily focus on driving scenario generation [Ni2025MaskGWMAG, Hassan2024GEMAG, li2025semi], planning [wang2023driving, liu2025qfft], and representation learning [Min2024DriveWorld4P, Yang_2024_CVPR, zeng2024driving]. For driving scenario generation, most prior works are built upon diffusion models, with the exception of GAIA-1 [Hu2023GAIA1AG] which incorporates a progressive next-token predictor and an additional diffusion image decoder. Recent DrivingGPT [Chen2024DrivingGPTUD] leverages existing vision generation LLM LlamaGen [sun2024llamagen] while simultaneously outputting predictions for future states and actions. However, such VQ-VAE based visual tokens lack semantic information, often leading to performance degradation in downstream visual understanding tasks [xie2024showo, liu2024rag, tan2025teaching]. In this work, we propose to directly activate the visual generation capabilities of existing multimodal large language models, enabling VLMs to act as world models and predict future frames.
@@ 3.1 Preliminary
End-to-end autonomous driving directly generates future trajectory from sensor data, convertible to vehicle control actions like acceleration and steering [Hwang2024EMMAEM]. given $N$ surround-view images $I_{t}=\{I_{t}^{1},I_{t}^{2},\ldots,I_{t}^{N}\}$ at timestep $t$, model $\mathcal{M}$ outputs a BEV trajectory $W_{t}=\{w_{t}^{1},w_{t}^{2},\ldots,w_{t}^{n}\}$, where each waypoint $w_{t}^{i}=(x_{t}^{i},y_{t}^{i})$. The process is formulated as:
@@
Recent works [wu2024janus, huang2024magicfight] unify multimodal understanding and vision generation in single LLM. While understanding aligns with standard LLMs, generation methods [lwm, huang2025m4v] typically use VQ-VAE [vqvae] to encode images into discrete tokens. First, the image tokenizer quantizes image pixels $x\in\mathbb{R}^{H\times W\times 3}$ into discrete tokens $q\in\mathcal{Q}^{h\times w}$, where $h=H/p$, $w=W/p$, $p$ is the downsampling factor, and $q(i,j)$ represents the index of the image codebook. These $h\cdot w$ tokens are arranged in raster order to train a Transformer [Vaswani2017AttentionIA]-based autoregressive model. During image generation, a general language modeling (LM) objective is adopted to autoregressively predict the next token, maximizing the likelihood of each image token:
@@ 3.2 Unified pre-training paradigm for visual generation and understanding
To enable unified pre-training, MLLMs require visual generation capabilities. As described in Section 3.1, existing methods (e.g. Lumina-mGPT [liu2024lumina-mgpt], the visual generation LLM used by Doe-1 [doe]) typically employ VQ-VAE to encode images into discrete tokens when extracting visual information. However, these tokens lack semantic information, which hurts downstream understanding performance [xie2024showo, zhou2025opening]. Moreover, current methods [wu2024janus, zhou2025foodsky] demand expensive training from scratch on massive billion-scale datasets without leveraging existing LLM knowledge.
@@
To effectively preserve the semantic understanding capabilities of the native MLLM during the pre-training stage, as shown in the left part of Figure 2, we follow previous methods [Wang2024OmniDriveAH, Hwang2024EMMAEM] by using a VQA task, which is crucial for autonomous vehicles to analyze complex driving scenarios. Given an image-text question-answer pair $(I,L)$, where $I$ represents the surround-view images of the current scene and $L$ denotes the instructional question, model $\mathcal{M}$ generates a corresponding answer $A$:
@@
Inspired by the world models in autonomous driving [kim2021drivegan, yang2024genad] that generate future frames to learn physical laws, after activating the visual generation capability, we also enable the VLA to predict future frames, thereby capturing the dynamic evolution of the world. Specifically, given an image-instruction pair $(I,L)$, the model predicts the next visual token of the future front-view frame through autoregressive generation:
@@
However, directly generating complete detailed future scenes may fail to adhere to physical laws [Yang2025VLIPPTP]. Therefore, during pre-training stage, we propose a progressive, easy-to-hard generation method, incorporating annotated data containing lane divider and 3D detection. Before generating visual tokens of future frames $Q_{f}$, we leverage the world knowledge of VLA to first reason about visual tokens of lane dividers $Q_{l}$, which serve as the skeleton of the road scene and define drivable areas to enforce static physical constraints. Subsequently, we reason about visual tokens of 3D bounding boxes $Q_{d}$, representing motion patterns of key objects to impose dynamic physical constraints. This progressive method sequence explicitly guides the model to infer structural layouts and geometric details of future scenes while enforcing physical laws. By leveraging these intermediate visual reasoning steps as context, the model learns to think visually about the dynamic evolution of scenes, ultimately enabling accurate future prediction:
@@ 3.3 Think visually with spatio-temporal CoT
Autonomous driving planning requires not only understanding the current scene but also envisioning potential future developments to achieve forward-looking comprehension. This thinking process should resemble physical world simulation and imagination rather than purely text symbolic logical deduction. Since our model has already learned physical constraints through the progressive generation during pre-training, and considering efficiency, we no longer separately generate lane dividers, 3D detection, and future frames, but instead integrate all these results into a single unified frame. As shown in the right part of Figure 2, here, VLA serves as a world model to generate a unified image frame predicting the future world state: Inspired by visual prompting engineering [redcircle] that draws red circles on images to guide model attention and by VLIPP [Yang2025VLIPPTP] first predicts future bounding boxes to introduce physical priors when generating future frames, we represent future world spatial relationships through future red lane dividers and 3D detection boxes on the predicted unified frames. These coarse-grained visual cues direct the model’s attention toward drivable areas and critical objects in future scenes while enforcing physically plausible constraints. Meanwhile, the temporal relationships are represented by the ordinary future frame, where the dynamic evolution of visual content intuitively characterizes temporal progression and the inherent laws of scene development. Subsequently, spatio-temporal CoT $Q_{CoT}$ serves as an intermediate reasoning step, allowing the VLA to function as an inverse dynamics model that plans trajectory based on current observations and future predictions: